UT Sep 07, 2019 - Restart Night #6


OSA: S. Allanson

ISp: C. Bolyard

ISA: O. Kuhn (Tucson)

SO: J Power (Tucson)

Mtn Manager: P. Hartley

Technical Expert: J. Hill (Tucson)

Plan for Tonight:
  • Bino-LUCIs to start
    • PFSAlign for SX
    • Twilight Flats - need clear skies
    • AGW-1 and AGW-2 Transform - high (2x2hrs - requires initial data, then followup to confirm transform)
      • LUCI-1: x-shift 4.84" y-shift -1.14" theta = 0.067
      • LUCI-2: x-shift 3.61" y-shift -2.28" theta = -0.260
      • LUCI-2 PCS config needs to be updated to factor in the theta
        • Make the change in the PCS, update, restart systems, then take another mask acquisition to verify that the theta ~ 0 - this is because we often get the sign backwards
        • Take the data
        • Once transform is completed, revert PCS change
    • non-sidereal check imaging - high (0.5hr) - need this to verify all facility instruments work OK in NS with TCS 2019A
    • M3 alignment DX (lower priority)
  • LBCs
    • Superfocus sequence -PA=0,180 high (1 hr)
    • Test Science OB
      • Sidereal imaging/dither tests (never completed this due to LBC-B issue)
      • NS Imaging - C. Veillet object 2nd half - need this to verify all facility instruments work OK in NS with TCS 2019A
    • LBC Subframing (Use test version of OT and check chip 4 subframing on sky)
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations - need clear skies
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations - need clear skies
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete) - need clear skies
  • MODS (if time or conditions permit)
    • non-sidereal guiding with longslit observations (object stays in slit without issue) - test of new PCS/GCS/TCS
  • CLOSED DOME:
    • Highest Priority is continued flexure testing Bino-LUCis
    • Revert to main page for additional closed dome tests - also can do other tests in parallel

Summary:
  • CRITICAL IT LUCI1 - MCE 7583. Ran LUCI2 only to begin.
  • AGW-2 Transform -initial data. /home/lbto/data/20190907/rfdg/transform_luci2.log
  • non-sidereal check imaging with dither- "guide loop didn't pause within timeout" IT 7845 FAIL
  • sidereal check imaging - PASS
  • LUCI Rotator center check
  • LBC Test science OB Non Sidereal with and without dithers - guiding failed periodically even when tracker images looked ok.
    Take 8 sec but did not appear to attempt 16, 32 or higher iterations.
  • LBC Test Science OB Sidereal
  • LBC SUPERFOC
  • Skyflats g,i, Uspec, z


  • IT 7638 Tripped Breakers CB1 & CB4 when trying to powerup LUCI1 MCE & ROE at the start of the night
  • Chris reset the breaker.
  • We were able to power on after waiting 20 minutes since the logout was not clean.
  • LUCI1 MCE not responding. Camera did not init. Checking LUCIFER MCU none of the motors are returning status. Chris could hear fans in the MCE. Called Dave. We cycled the LUCI1 software no change, tryed pinging the various LUCI1 IP's nothing amiss. Chris talked to James and he will take a peak to see if there are any LED's on the MCE. No LEDS, no power. LUCI 1 is kaput for now.

OSA: S Allanson Observers: J Power and O Kuhn Technical Expert: J Hill


02:15 Opened. Going to focus on collecting transform data for LUCI2 only. Will check RZP and update it in PCSConfiguration first. (NB for MODS, if stars need to rotate CW to be in boxes, then subtract dtheta from RZP, but there may be a sign flip for Bent Gregorian stations).

02:52 MOS alignment. Using RTD on 64-bit machines as we did a couple nights ago, but getting different behavior. The box overlay does not appear well aligned with the mask image, after the mask-to-detector transformation was done, and not all of the stars are used in the first cut - stars 30 and 33.. The box positions were perfect a few days ago, but it appears to be slightly off today. Jenny calling Igor. Igor suggested clicking "Subwindow Requery using FITS keywords". Doing that, but it seems to be stuck at "waiting response..." (it's been there for minutes). Igor says nothing has changed in this part of the software that predicts reference star positions.

03:15 Clouds coming in. MOS alignment can be done, but stars 30 and 33 did not show up. 30 is in the bad channel but a bright star.

 MOS align based on 4,5 and 7 (mask): 
dx = 3.22727, 
dy = -2.44855, 
dtheta = -0.24850, 
error = 0.05914. 

Consistent with 20190905 values for this mask with LUCI-2 (from 20190905 log: x-shift 3.61" y-shift -2.28" theta = -0.260).

03:34 Sending offset. luci2.20190907.0008 is verification image. Looks good.

03:40.

cp PCSInstrument.conf PCSInstrument.20190907.conf

RIGHTZEROPOINT 176.645 currently. 

03:56 Edited PCSInstrument.conf to have RIGHTZEROPOINT 176.895. Steve restarted PCS.

04:00 Jenny resending the preset and taking MOS alignment images to check RZP.

04:08 No change (dtheta was -0.266). Olga updated PCSInstrument.conf file in wrong location (did so in /lbt/tcs). In this set of alignment images 9,10 and 11, the stars 30 and 33 showed up. All stars showed up and errors were very small once one star (34) that had a bright companion and had been mis-ID'd by software was reselected manually.

04:09: Correct path is: /lbt/data/config/tcs/PCS on tcs1. In that PCSInstrument.conf file, RIGHTZEROPOINT was 177.05. Copied the current version to PCSInstrument.20190907.conf and then, in PCSInstrument.conf, I set RIGHTZEROPOINT to 177.30. Steve restarted PCS.

04:10: Jenny resending the preset and taking MOS alignment images with RZP 177.30. These will be 12, 13 and 14. Alignment was good. The offsets indicated are:

MOS align based on 12,13 and 14 and using RZP 177.30 are:
dx =  2.236465
dy = -2.91095
dtheta  = -0.01033
error = 0.05590 

04:30 Resending once more: 15, 16 and 17.

 MODS align based on 15,16 and 17 and still using RZP 177.30:
dx = 2.43753
dy = -2.93084
dtheta = 0.00956
error = 0.05577 

We're going to stick with this RZP and move on to collecting transform data, using Stone O (Stone L is setting) and obs2:/home/lbto/idl/wfsc/transform/lists/Stone_Opm10.list

04:35 Resending preset because we are seeing the guide star overshoot the hotspot. We saw this on about 3 of the last 3 preset, but not this one. Small number stats? (IT xxxx).

04:40 Prep for transform.
  • Steve tuning pointing and collimating at Stone O.
  • Jenny moving in N30_FieldStop mask and setting filters to clear + Ks.
  • We had some question about whether to use the N30 or N3.75? We are assuming we'll take data with the N3.75 but if the AFC spots are used for reference then either camera would be OK.
  • Jenny aligned the Field Stop - it was a bit off.

04:55 Starting transform_collect using StoneO _pm10.list (on-axis star is 1):
luci2 GS
24 1
24 and 25 are with Ks in place, but the preset wavelength is at 1.65 (look in IIFGUI Data).

Jenny switched to H and is checking the exposure time. Setting the exposure time to 4.

05:02 Running transform_collect. Images are luci2.20190907.00xx.fits
xx GS
27 1
28 background
05:06 Error file not found and transform_collect crashed. We saw a message about wfs info file was not found in daily images.

05:12 transform_collect, exp=4.0

05:15. Guide probe could not reach star. Stone_Opm10.list was a list for MODS!

cp Stone_Opm10.list Stone_O180pm10_luci.list

and add 180 deg to all PAs

05:15 transform_collect, exp=4.0 now with Stone_O180pm0_luci.list. Seeing is good. H-band image has 0.54" FWHM.
31 on-axis star 1 PA=180  
32 background  
33 11  
34 15  
35 24  
36 23  
37 28  
38 26  
39 33  
40 32  
41 12  
42 10  
43 back on-axis PA = 190, background image
44 11 looks like there is a shift in star position
45 15  
46 24  
47 23  
48 28  
49 26  
50 33  
51 32  
52 12  
53 10  
    collimating on-axis at PA=170
54 on-axis PA=170, background image
55 11
56 15
    collimating on-axis
57 24  
58 23 seeing ~0.5-0.75"
59 28  
60 26  
61 33  
62 32  
63 12  
64 10  
65 1  
Done. Log is /home/lbto/data/20190907/rfdg/transform_luci2.log

06:08 Going on to NS imaging. We'll try Triton (slow) and 2014 A (faster). Triton is close to an extremely bright.

Reverted PCSInstrument.conf to use old RZP 177.05 and Steve is restarting PCS.

cp PCSInstrument.conf PCSInstrument.20190907_NEW_LUCI_RZP.conf
edit PCSInstrument.conf to revert the RIGHTZEROPOINT 177.05
Steve restarting PCS

06:27 Pointing Check. Looks good.

06:30 Sending imaging script on 2014 AF51.

06:35. Error on first offset. Guide loop didn't pause within timeout.

Sat Sep 7 06:37:46.500 2019 right Offset alert for LUCIFER: guide loop didn't pause within timeout
Sat Sep 7 06:37:46.552 2019 right Offset command failed for LUCIFER, Sound=offsetFailed.wav

06:36 Sending it again. Same error.

06:38 Sending same script but without NS override to see whether the problem is with NS or general. This time, the guide loop paused.

06:43 Steve restarted GCS. Jenny resending the NS preset. The guide loop is now pausing for the offsets. Looks good. Object seen. The set of images is 0066-0071.

06:57, however, we got another one of the "guide loop didn't pause within timeout"

Sat Sep 7 06:57:23.313 2019 right Offset alert for LUCIFER: guide loop didn't pause within timeout
Sat Sep 7 06:57:23.342 2019 right Offset command failed for LUCIFER, Sound=offsetFailed.wav

07:00 The subtracted image: 68 minus 69 is displayed in a weird way by Aladin. This is still for the NS object. Earlier subtractions were displayed normally, it was only on this pair that there was a problem - something odd about the WCS in the FITS headers? or something about the RTD/Aladin?

07:04 Another error with the "guide loop didn't pause within timeout". John looking into it. Timeout is 6.0 sec and actual pause guiding, in one instance at least, took 6.x seconds.

07:10 Running sidereal dithering script. M52, 9-pt dither, PA=0. This will be the sequence starting with index number 0072-0076. No offset problems.

07:30 Next up, rotator center check with LUCI2 on M52, using N30_FieldStop. The script will take sky, then 4 PAs.
78 sky offset 300" 300"
79 PA=0

78 and 79 were for an active preset. Track presets now: to PA=0, PA=180, PA=xx, PA=xx.
81 sky @ 300" 300"  
82 PA=0  
83 PA=180 looks pretty poorly collimated
84 PA=90 7pix, but DIMM=1.68", we were used to the <0.6" images!
85 PA=270
86 sky
07:50 Done with LUCI2 work. We took data to measure the transform for LUCI2/DX, measure the rotator center, and test sidereal & NS guiding + dithering.

07:56 Reconfiguring to LBCs.

08:06 Taking a few darks with LUCI for the transform data while reconfiguring. Blind filters, N30 fieldstop in. DIT 4s NDIT1 NORMAL luci2.20190907.0087-91

08:09 LBC connect to LBT

08:14 Reconfigured for LBC's. First up, checking non sidereal with LBC's. Sending for copointing near 2018 YH.

08:17 Running do hybrid. Dohybrids models match reality today.

08:26 FPIA converged. Copointing. Noticed a strip on the blue on the right of chip 2. Also on red but no structure This was not seen in RB science but noted in the lbcrangebal image. It was the overscan. Thrown off by a gradient at the base of the overscan.

8:50 blue guiding failed 084517

8:54 Moving on to 2015 AH51 non sidereal, copointing to start.

8:58 DOFPIA. Clear skies. NS target at 40 deg el and setting, DIMM 1.15". During DOFPIA seeing blowing up, 2.16", but 4.9pixels on blue for copointing image. Red are not as well collimated but about the same FWHM ~4.8pixels

9:15 Hijacked Preset to 2014 AF51. Starting with stare script. 5x3 min exp.

9:32 Blue failed 092944 failed. We aren't dithering and seeing between 1.25-1.75". Did not appear to try to go on above 8seconds. Should have tried 16, 32sec when 8 failed but did not appear to do so.

9:35 Taking sidereal image of field for IQ purposes. Red looks a bit rough reading a FWHM of 1.6-2.5 on trackers, blue a bit better.

9:39 Sending Nonsidereal LBC dither script

9:51 Moving to Little Dumbell copointing. Lots of stars around so that if the IQ is ok we can go a head with a superfoc sequence.

9:56 Running dofpia. DIMM seeing 1.2-1.5" and copointing x2. Copointing corrections very small. From Copointing:Blue 3.6pixels, Red 3.6pixels.

10:16 IRAF issues, take down and bring back up. Rerunning DOFPIA again

10:20 /home/lbto/Calib_OBs/SUPERFOC, Running Superfoc RB_rVsf7 at PA0

10:27 RB_rVsf7 Superfoc PA 180

10:34 RB_rVsf7 Superfoc PA 0

10:41 Running dofpia again

10:45 RB_rVsf7 Superfoc PA 0

10:53 RB_rVsf7 Superfoc PA 180

10:59 DOFPIA sequence on little dumbell

11:05 Little Dumbell (M76) Science OB Cycling filters and dithering. Seeing 1.24" on DIMM. "Too many iterations, unable to compute sky" on FPIA happens on crowded fields. Red 4.35 pixels, Blue 4.4 pixels. Nice and round

11:29 Red image 112807 bad image 112821 too. Aberrations.

11:31 DOFPIA Quite a large focus correction, particularly for blue. Focused on field so some of the nebula in there but no issues with "too many iterations, unable to compute sky"

11:36 Running Science OB to complete the last 2 dithers on Little Dumbell. 0.9 & 0.7" FWHM in trackers

11:56 Running a dither test with 4 point dither single filter. Note that the IIF says we are sending Async preset and Offsets, while we are clearly sending sync. Olga found an email from Kellee that explained that all Binocular LBCR/LBCB Presets and Offsets are sync unless explicitly stated.

12:05 The IQ looks nice so rerunning the 4 pt dither science OB again. Sun is sneaking up on us.

12:10 Slewing to BlankField03 +31 for skyflats

g , i - i started too bright, g good

PA = 0
lbcb.20190907.121457.fits SkyFlat g-SLOAN 15.270 16K
lbcb.20190907.121605.fits SkyFlat g-SLOAN 15.282 19K
lbcb.20190907.121712.fits SkyFlat g-SLOAN 15.270 24K
lbcb.20190907.121958.fits SkyFlat g-SLOAN 12.782 33K
lbcb.20190907.122057.fits SkyFlat g-SLOAN 12.771 40K
lbcb.20190907.122226.fits SkyFlat g-SLOAN 2.782 11K
lbcb.20190907.122314.fits SkyFlat g-SLOAN 2.770 13.3K

lbcr.20190907.121452.fits SkyFlat i-SLOAN 26.918 43K
lbcr.20190907.121601.fits SkyFlat i-SLOAN 26.910 54K
lbcr.20190907.121708.fits SkyFlat i-SLOAN 26.918 60K
lbcr.20190907.121950.fits SkyFlat i-SLOAN 5.235 21K
lbcr.20190907.122052.fits SkyFlat i-SLOAN 5.244 25.3K
lbcr.20190907.122224.fits SkyFlat i-SLOAN 1.235 7.5:
lbcr.20190907.122310.fits SkyFlat i-SLOAN 1.244 8.6K
lbcr.20190907.122503.fits SkyFlat i-SLOAN 2.014 21K

g, i PA180
lbcb.20190907.122506.fits SkyFlat g-SLOAN 1.284 8.5K
lbcb.20190907.122551.fits SkyFlat g-SLOAN 1.267 10K
lbcb.20190907.122634.fits SkyFlat g-SLOAN 1.284 12K
lbcb.20190907.122723.fits SkyFlat g-SLOAN 1.267 15K
lbcb.20190907.122807.fits SkyFlat g-SLOAN 1.284 18K

lbcr.20190907.122545.fits SkyFlat i-SLOAN 2.023 25K
lbcr.20190907.122631.fits SkyFlat i-SLOAN 2.014 29K
lbcr.20190907.122719.fits SkyFlat i-SLOAN 2.022 34K
lbcr.20190907.122803.fits SkyFlat i-SLOAN 2.014 40K

Moving to PA180

USpec and z at PA180

lbcr.20190907.122938.fits SkyFlat z-SLOAN 1.244 18K
lbcr.20190907.123031.fits SkyFlat z-SLOAN 1.236 22K
lbcr.20190907.123121.fits SkyFlat z-SLOAN 1.244 27K
lbcr.20190907.123211.fits SkyFlat z-SLOAN 1.236 33K
lbcr.20190907.123301.fits SkyFlat z-SLOAN 1.244 40K

lbcb.20190907.122948.fits SkyFlat SDT_Uspe 3.815 13K
lbcb.20190907.123036.fits SkyFlat SDT_Uspe 3.831 16K
lbcb.20190907.123125.fits SkyFlat SDT_Uspe 3.815 20K
lbcb.20190907.123219.fits SkyFlat SDT_Uspe 3.831 25K
lbcb.20190907.123304.fits SkyFlat SDT_Uspe 3.816 31K

USpec and z at PA0

lbcb.20190907.123521.fits SkyFlat SDT_Uspe 2.056 28K
lbcb.20190907.123606.fits SkyFlat SDT_Uspe 2.045 34K
lbcb.20190907.123648.fits SkyFlat SDT_Uspe 2.057 40K

lbcr.20190907.123516.fits SkyFlat z-SLOAN 0.735 38K
lbcr.20190907.123600.fits SkyFlat z-SLOAN 0.743 45K
lbcr.20190907.123645.fits SkyFlat z-SLOAN 0.735 54K

Olga noted a dark out of focus feature on chip 2 in the flats. Upon closer inspection this appears in multiple filters and has been on science images through the night. It is particularly visible in the g image and Uspec image. It apears like a doughnut on chip 2. IT 7846.

Took a series of 25 biases with the tel @ zenith, and then another set of 25 biases with tel @ horizon to look for orientation effects in the LBCR horizontal banding. First glance, looks like the banding is present in zen biases and not in horizon ones.

-- BarryRothberg - 06 Sep 2019
Topic revision: r8 - 09 Sep 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback